Brian and Daryl went shopping. both decided to spend the 532 pesos they have saved from their allowance. Daryl spent 1/2 of her money. After Brian spent 82 pesos,they had the same amount of money left. How much money did Brian have at first?
Answer by Boreal(15235)

Brian had x
Daryl had 532-x
Daryl spent (532-x)/2
Brian spent 82
B now has x-82
Daryl now has (532-x)/2
those are equal
multiply both sides by 2
2(x-82)=532-x
2x-164=532-x
3x=696
x=232 for Brian
Daryl has 532-232=300
Brian spent 82 and has 150
Daryl has half of 300 or 150